# A Dense Subframe-based SLAM Framework with Side-scan Sonar

### v1

@@ Abstract
Side-scan sonar (SSS) is a lightweight acoustic sensor that is commonly deployed on autonomous underwater vehicles (AUVs) to provide high-resolution seafloor images. However, leveraging side-scan images for simultaneous localization and mapping (SLAM) presents a notable challenge, primarily due to the difficulty of establishing sufficient amount of accurate correspondences between these images. To address this, we introduce a novel subframe-based dense SLAM framework utilizing side-scan sonar data, enabling effective dense matching in overlapping regions of paired side-scan images. With each image being evenly divided into subframes, we propose a robust estimation pipeline to estimate the relative pose between each paired subframes, by using a good inlier set identified from dense correspondences. These relative poses are then integrated as edge constraints in a factor graph to optimize the AUV pose trajectory.
The proposed framework is evaluated on three real datasets collected by a Hugin AUV. Among one of them includes manually-annotated keypoint correspondences as ground truth and is used for evaluation of pose trajectory. We also present a feasible way of evaluating mapping quality against multi-beam echosounder (MBES) data without the influence of pose. Experimental results demonstrate that our approach effectively mitigates drift from the dead-reckoning (DR) system and enables quasi-dense bathymetry reconstruction. An open-source implementation of this work is available111TBA.
@@
Many commercial autonomous underwater vehicles (AUVs) rely on the dead-reckoning (DR) system for underwater navigation, which is subject to unbounded errors coming from the accumulation of sensor uncertainties. The common solutions to mitigate such errors involve either integrating global referencing systems (i.e., GPS) that requires periodic resurfacing of the vehicle, or deploying pre-installed acoustic ranging systems such as long/short/ultrashort baseline setups, which could be cumbersome to deploy and confined to limited operational ranges. An alternative solution is to incorporate sensor measurements of the environment to reduce the dead-reckoning drift using a simultaneous localization and mapping (SLAM) method [1].
@@
For AUVs surveying undersea, three types of sonar sensors are frequently-used: side-scan sonar (SSS), useful for producing high resolution images of a large swath of the seafloor; multi-beam echosounder (MBES), useful for producing the D geometry of the seafloor directly; and forward-looking sonar (FLS), useful for capturing wide-angle, fan-shaped images at consecutive times with large amount of overlap in between. MBES provides D data of the seabed as point clouds that are relatively easy to interpret for underwater SLAM applications [2][3]. The main drawbacks of using MBES for SLAM, however, are its relatively narrow coverage resulting in limited data overlaps between surveying paths, and the difficulty to recognize patterns in terrains that are geometrically featureless, e.g., relatively flat or with gentle slopes. The typical range of FLS is from meters to over a few tens of meters, making it suitable for robust SLAM in small-scale environments [4][5][6]. Nevertheless, the utilization of forward-looking sonar comes with unique challenges, including low image texture and high signal-to-noise ratio (SNR). In contrast, SSS is generally used to map seafloor of hundreds and thousands of meters in large-scale environments. Compared to MBES, SSS offers higher resolution that allows for precise AUV localization, as well as larger coverage that allows for localization and mapping using large areas of the seafloor.
@@
However, the problem of SLAM with SSS images is far from being solved, mainly due to the challenge of registering SSS images to one another and the lack of D information. In particular, raw SSS images are geometrically distorted and unevenly ensonified, making the appearance of same region of seabed vary in SSS images when being observed from different positions. To address this, we first apply a canonical transformation to the raw SSS images to reduce such distortions and redistribute the image pixels to be approximately equal size patches of seabed. Then we propose an effective dense (pixel-wise) matching method based on [7], combining both geometric and appearance constraints of side-scan image, to find dense correspondences between overlapping images.
@@
The inability of measuring D data makes the estimate of AUV pose underdetermined using single-ping measurement, even with accurate associated correspondence [8]. To mitigate such underdetermination, we propose to evenly divide each SSS image into subframes and estimate the relative pose of the centre pings between paired subframes using multiple ping measurements from dense correspondences as constraints. To be robust against noise and outliers in the dense correspondences, we integrate the pose estimation into a RanSaC (Random Sample Consensus) [9] pipeline, using both measurement cost and optimization cost as iteration signal.
@@
With the above considerations, we propose to estimate the AUV poses and landmarks of paired correspondences as a subframe-based graph SLAM problem solved in two steps. First, we model the selected pixel correspondences in each paired of associated subframes as D measurements, which are formulated together with dead-reckoning constraints as a least-squares minimization problem in a robust estimation framework, to deliver an accurate relative pose constraint between the two associated subframes. In the second step, all of such estimated poses are considered as loop-closure constraints in a pose graph for global optimization that refines the entire AUV pose trajectory. Compared to sparse ping/feature-based approaches [8][10], our proposed method is able to perform more accurate AUV pose estimation, as well as generate quasi-dense bathymetry point cloud using the dense correspondences and optimized pose trajectory.
@@ II-A SLAM with Side-scan Sonar
Research on SLAM with SSS has drawn considerable attention within the community in the past two decades. Among the earlier works, [11][12][13] combine a stochastic map with Rauch-Tung-Striebel (RTS) filter to estimate AUV’s location, where the stocahstic map is formulated as extended Kalman filter (EKF), with landmarks manually extracted from the side-scan images. Later in [14], the authors propose to fuse acoustic ranging and side-scan sonar measurements in pose graph SLAM state estimation framework that avoids inconsistent solution caused by information lost during linearization [15], to achieve accurate AUV trajectory estimation. Similarly, [16] proves the capabilities of using SSS images in graph-based iSAM [17] algorithm to produce corrected sensor trajectories for image mosaicing. Issartel et al. [18] take a further step to incorporate switchable observation constraints in a factor graph to address false data association issue and guarantee robust solution. The above solutions prove the feasibility of using SSS information to help better localize AUVs, however, they rely on manual annotation of landmarks in SSS images, which goes against full autonomy in AUVs.
There are but few approaches attempting to concurrently address the problems of both SSS image association and SLAM. [10][19] proposes to use a boosted cascade of Haar-like features to perform automatic feature detection in SSS images. The correspondences are found by means of joint compatibility branch and bound (JCBB) algorithm [20], and used to update pose states in an EKF SLAM framework. By applying a similar framework, [21] combines local thresholds and template-based detectors with strict heuristics to avoid false associations. The integration of JCBB in EKF is arguably well-fitted to solve the complete SSS SLAM problem with fewer landmark measurements, but it suffers from inconsistent performance and offers lower accuracy when comparing to graph-based SLAM solution [22].
@@ II-B Feature Detection and Matching of Side-scan Image
Some efforts have been made to exploit feature extraction and matching in side-scan images, mainly for estimating the relative pose of AUV through image registration. [23] is one of the earliest works that try to apply feature-based registration method using SIFT [24] for side-scan sonar images. Later in [25], the authors extensively compares the performance of off-the-shelf feature detection techniques in matching side-scan sonar image tiles, and integrate some of the robust solutions into an image registration framework for their AUV route following system [26]. Lately, the fusing of visual and geometric information to accomplish data association becomes a trending solution. For instance, [27] uses elevation gradients as additional features to assist in associations between landmarks. In [28], the authors propose to construct features with D location and feature intensity, and associate between them through a modified version of iterative closest point (ICP) [29] scan matching method. While interesting conclusions are drawn in the above works, none has integrated an applicable solution into a full SLAM framework.
@@ II-C Dense Matching of Optical Image
Large displacement optic-flow estimation is a task to establish pixel-wise correspondences between optical images with large parallax, and finding dense correspondences between parallel side-scan images is closely related to this task. While classic variational approaches [30][31] have been tried by integrating sparse descriptor matching in a cross-to-fine scheme to mitigate the large displacement issue, they typically assume local smoothness and that motion discontinuities only happen at the image edges [32], which do not apply to side-scan images, since there are discontinuities at the nadir of each side-scan image and when the trajectories of forming parallel side-scan images are close to each other, see Fig. 2.
In contrast, methods [33][34][35][36] based on nearest neighbor fields (NNF) suit better to our problem as they are able to perform efficient global search for best match, directly on full image resolution. Nevertheless, NNF-based methods usually generate many outliers that are difficult to identify, due to lack of regularization. In this paper, we propose an effective dense matching method for side-scan image, by carefully modifying on a NNF-based seminal work called PatchMatch [7] to combine both dead-reckoning and imaging information to search for dense correspondences. To mitigate the influence of outliers to pose estimation, a robust RanSaC-based pose estimation pipeline is introduced.
@@ III-A1 Canonical Transformation
Data association across different SSS images is challenging, as the same physical area on the seabed will appear differently in the SSS images collected from different distances and angles. Furthermore, an SSS image will be distorted compared to an orthographic projection of the ensonified region and will vary in intensity as a function of the sonar position. To reduce such distortions, we apply a method [37] capable of transforming SSS images from different survey lines into a canonical representation, which includes two steps: intensity correction and sensor independent slant range correction.
Intensity correction is performed to normalize the intensities across images based on the ’Lambertian’ model [38][39], with the SSS backscatter intensity being proportional to the cosine square [40] of the incidence angle. The slant range correction adjusts for the varying projection of side-scan pixels on the assumed horizontal seafloor. An adjustment of sine of the incidence angle is performed to be a distance along the horizontal, so that the pixels can be of fixed size in terms of horizontal range rather than slant range. More details can be found in [37].
@@ III-A2 Geo-referenced Image
To approximately check whether any two of the SSS images are overlapped, we utilize the dead-reckoning data to make each pixel in SSS image geo-referenced, i.e., approximate the location of each pixel of an SSS image in the global reference coordinate, following the method in [26]. This is also used to narrow the searching area when performing dense matching in data association.
@@ III-B Data Association
Once an overlapping area is found between a new SSS image and any image in the database, a data association process is performed to find dense pixel correspondences between the overlapping images. In this work, we propose to accomplish this by formulating it as a - (NNF) estimation problem, and solving via computing the approximate NNF based on our modified version of PatchMatch [7], which is carefully tailored for side-scan images.
Specifically, define a nearest-neighbor field over all possible patch coordinates (locations of patch centers) in an side-scan image , so that given a patch coordinate in image , its corresponding nearest neighbor in side-scan image with overlap can be denoted as,
The values of are referred as and stored as a matrix with the same dimension of the image size of . Following the pipeline in [7], the estimation of has three main steps: initialization, random search and propagation, see Algorithm 1.
@@ III-B2 Random Search
Based on the initialized offset at each position, a random search is performed around it within a , which can be determined by prior knowledge of navigation drift and imaging resolution. Then given a random selected position, compare the patch distance against its current position, if it becomes smaller, then replace the current with new selected position. The distance function we use in this paper is Zero-mean Normalized Cross-Correlation (ZNCC) distance, which is invariant to linear brightness and contrast variations that often appear close to the nadir and the sides in side-scan images.
@@ III-B3 Propagation
Assuming that the patch offsets are likely to be the same locally, the propagation step aims to improve the values of using the known surrounding offsets. Specifically, given being the patch distance between the patch at in image and the patch at in , we take the new value for to be the minimum patch distance within the neighbors:
The effect is that if has a correct mapping and is in a coherent region, then all of neighbors around will be filled with the correct mapping. However, this could be an issue for side-scan images, which have disconnection in centre column after removing the nadir via canonical transformation, and the noisy areas on the top and bottom due to the AUV making turns, as well as sensor artifacts. To address this, we mask out these areas and exclude them from the matching process.
@@ III-B4 Improving Data Association using Optimized Poses
The presets of patch size and max offset cannot be too high as that would in general lower the patch matching accuracy, which makes it difficult to handle large offsets/drifts in cases where the paired images are distant from each other. To address this, we propose to improve the dense matching accuracy in distant paired images with optimized poses in an iterative fashion (see Fig. 3), which in turn helps to improve the pose accuracy. Here a new parameter representing the iteration number of this process is introduced. Detailed discussion can be found in Section IV-C2.
@@ III-C Subframe Pose Estimation
Now we show how each pair of the pixel correspondences is used to construct a measurement model. Based on that, we propose a robust pose estimation pipeline, which is able to estimate the relative pose between a pair of subframes of SSS images, by iteratively identifying a set of good quality inliers from dense correspondences using RanSaC. In the following, we assume that each side-scan image has been evenly divided into multiple subframes and they are used as unit image for estimation instead of the whole one, see Fig 4 left.
@@ III-C2 Relative Pose Estimation
Given a set of pixel correspondences between two overlapped subframes, the observed D landmarks and centre ping poses of the subframes can be estimated via minimizing the minus log of least squares cost, assuming the measurement noises as Gaussian:
with the first term being the side-scan measurement cost discussed above, the second term the odometry measurement cost and the third term a prior cost. Specifically, is the relative pose between the centre pings of the subframes, and is the measurement obtained from dead-reckoning data. The odometry term is crucial as it helps to relieve the underconstrained issue in side-scan measurement and ensure convergence to a desired minimum. is a prior model that sets one of the poses (e.g., here) fixed and only adjust for the other, which is reasonable as we aims to estimate the relative pose between them rather than estimate both.
is the covariance matrix for each side-scan bin measurement. The variance in the range measurement is dominated by the discretization of the range in the side-scan, varying from - depending on the resolution of the image. For the second term restricting the landmark to lie in the plane, the variance should grow as the square of the horizontal distance from the sonar. Since the sonar is mostly used at shallow grazing angles, we can instead use the range so that:
where is the beam width in radians. is the odometry covariance and is set proportional to the distance between the two dead-reckoning poses. A factor graph describing Eq. 9 is illustrated in Fig. 4 bottom right.
@@
The core idea of this pipeline is the consideration of both measurement cost ( and ) in the unsampled data, and in the sampled data. The former measures how well the hypothesis model fits to the unsampled data, while the latter indicates how well the optimization converges using the sampled data. We found experimentally that only using the measurement cost to determine the best model could be biased in many cases, while combining with optimization cost helps mitigate this issue by avoiding bad convergence. Therefore, we only update the current best model when both the costs are decreased.

### v2

@@ Abstract
Side-scan sonar (SSS) is a lightweight acoustic sensor commonly deployed on autonomous underwater vehicles (AUVs) to provide high-resolution seafloor images. However, leveraging side-scan images for simultaneous localization and mapping (SLAM) presents a notable challenge, primarily due to the difficulty of establishing sufficient amount of accurate correspondences between these images. To address this, we introduce a novel subframe-based dense SLAM framework utilizing side-scan sonar data, enabling effective dense matching in overlapping regions of paired side-scan images. With each image being evenly divided into subframes, we propose a robust estimation pipeline to estimate the relative pose between each paired subframes using a good inlier set identified from dense correspondences. These relative poses are then integrated as edge constraints in a factor graph to optimize the AUV pose trajectory.
The proposed framework is evaluated on three real datasets collected by a Hugin AUV. One of these datasets contains manually-annotated keypoint correspondences as ground truth and is used for evaluation of pose trajectory. We also present a feasible way of evaluating mapping quality against multi-beam echosounder (MBES) data without the influence of pose. Experimental results demonstrate that our approach effectively mitigates drift from the dead-reckoning (DR) system and enables quasi-dense bathymetry reconstruction. An open-source implementation of this work is available111https://github.com/halajun/acoustic_slam.
@@
Many commercial autonomous underwater vehicles (AUVs) rely on the dead-reckoning (DR) system for underwater navigation, which is subject to unbounded errors accumulated from the sensor uncertainties. The common solutions to mitigate such errors involve either integrating global referencing systems (i.e., GPS) that requires periodic resurfacing of the vehicle, or deploying pre-installed acoustic ranging systems such as long/short/ultrashort baseline setups, which could be cumbersome to deploy and confined to limited operational ranges. An alternative solution is to incorporate sensor measurements of the environment to reduce the dead-reckoning drift using a simultaneous localization and mapping (SLAM) method [1].
@@
For AUVs surveying undersea, three types of sonar sensors are frequently used: side-scan sonar (SSS), useful for producing high-resolution images of a large swath of the seafloor; multi-beam echosounder (MBES), useful for producing the D geometry of the seafloor directly; and forward-looking sonar (FLS), useful for capturing wide-angle, fan-shaped images at consecutive times with large amount of overlap in between. MBES provides D data of the seabed as point clouds that are relatively easy to interpret for underwater SLAM applications [2][3]. The main drawbacks of using MBES for SLAM, however, are its relatively narrow coverage resulting in limited data overlaps between surveying paths, and the difficulty to recognize patterns in terrains that are geometrically featureless, e.g., relatively flat or with gentle slopes. The typical range of FLS is from meters to over a few tens of meters, making it suitable for robust SLAM in small-scale environments [4][5][6]. Nevertheless, the utilization of forward-looking sonar comes with unique challenges, including low image texture and high signal-to-noise ratio (SNR). In contrast, SSS is generally used to map seafloor of hundreds and thousands of meters in large-scale environments. Compared to MBES, SSS offers higher resolution that allows for precise AUV localization, as well as larger coverage that allows for localization and mapping using large areas of the seafloor.
@@
However, the problem of SLAM with SSS images is far from being solved, mainly due to the challenge of registering SSS images to one another and the lack of D information. In particular, raw SSS images are geometrically distorted and unevenly ensonified. As a result, the appearance of same region of seabed can vary in SSS images when observed from different positions. To address this, we first apply a canonical transformation to the raw SSS images to reduce such distortions and redistribute the image pixels to be approximately equal size patches of the seabed. Then, we propose an effective dense (pixel-wise) matching method based on [7], combining both geometric and appearance constraints of side-scan image, to find dense correspondences between overlapping images.
@@
The inability of measuring D data makes the estimate of AUV pose underdetermined using single-ping measurement, even at the presence of accurately associated correspondence [8]. To mitigate such underdetermination, we propose to evenly divide each SSS image into subframes and estimate the relative pose of the centre pings between paired subframes using multiple ping measurements from dense correspondences as constraints. To be robust against noise and outliers in the dense correspondences, we integrate the pose estimation into a RanSaC (Random Sample Consensus) [9] pipeline, using both measurement cost and optimization cost as iteration signal.
@@
With the above considerations, we propose to estimate the AUV poses and landmarks of paired correspondences as a subframe-based graph SLAM problem solved in two steps. First, we model the selected pixel correspondences in each pair of associated subframes as D measurements. These measurements are formulated together with dead-reckoning constraints as a least-squares minimization problem in a robust estimation framework, to deliver an accurate relative pose constraint between the two associated subframes. In the second step, all of the estimated poses are considered as loop-closure constraints in a pose graph for global optimization that refines the entire AUV pose trajectory. Compared to sparse ping/feature-based approaches [8][10], our proposed method is able to perform more accurate AUV pose estimation, as well as generate quasi-dense bathymetry point cloud using the dense correspondences and optimized pose trajectory.
@@ II-A SLAM with Side-scan Sonar
Research on SLAM with SSS has drawn considerable attention within the community in the past two decades. Among the earlier works, [11][12][13] combine a stochastic map with Rauch-Tung-Striebel (RTS) filter to estimate AUV’s location, where the stocahstic map is formulated as extended Kalman filter (EKF), with landmarks manually extracted from the side-scan images. Later in [14], the authors propose to fuse acoustic ranging and side-scan sonar measurements in pose graph SLAM state estimation framework that avoids inconsistent solution caused by information lost during linearization [15], to achieve accurate AUV trajectory estimation. Similarly, [16] proves the capabilities of using SSS images in graph-based iSAM [17] algorithm to produce corrected sensor trajectories for image mosaicing. Issartel et al. [18] take a further step to incorporate switchable observation constraints in a factor graph to address false data association issue and guarantee robust solution. The above solutions prove the feasibility of using SSS information to help better localize AUVs. However, they rely on manual annotation of landmarks in SSS images, which goes against full autonomy in AUVs.
There are but few approaches attempting to concurrently address the problems of both SSS image association and SLAM. [10][19] proposes to use a boosted cascade of Haar-like features to perform automatic feature detection in SSS images. The correspondences are found by means of joint compatibility branch and bound (JCBB) algorithm [20], and used to update pose states in an EKF SLAM framework. By applying a similar framework, [21] combines local thresholds and template-based detectors with strict heuristics to avoid false associations. The integration of JCBB in EKF is arguably well-fitted to solve the complete SSS SLAM problem with few landmark measurements, but it suffers from inconsistent performance and offers lower accuracy when comparing to graph-based SLAM solution [22].
@@ II-B Feature Detection and Matching of Side-scan Image
Some efforts have been made to exploit feature extraction and matching in side-scan images, mainly for estimating the relative pose of AUV through image registration. [23] is one of the earliest works that try to apply feature-based registration method using SIFT [24] for side-scan sonar images. Later in [25], the authors extensively compare the performance of off-the-shelf feature detection techniques in matching side-scan sonar image tiles, and integrate some of the robust solutions into an image registration framework for their AUV route following system [26]. Lately, the fusing of visual and geometric information to accomplish data association has become a trending solution. For instance, [27] uses elevation gradients as additional features to assist with associations between landmarks. In [28], the authors propose to construct features with D location and feature intensity, and associate between them through a modified version of iterative closest point (ICP) [29] scan matching method. While interesting conclusions are drawn in the above works, none has integrated an applicable solution into a full SLAM framework.
@@ II-C Dense Matching of Optical Image
Large displacement optic-flow estimation is a task to establish pixel-wise correspondences between optical images with large parallax, and finding dense correspondences between parallel side-scan images is closely related to this task. While classic variational approaches [30][31] have been tried by integrating sparse descriptor matching in a cross-to-fine scheme to mitigate the large displacement issue, they typically assume local smoothness and that motion discontinuities only happen at the image edges [32]. These assumptions do not apply to side-scan images, since there are discontinuities at the nadir of each side-scan image and when the trajectories of forming parallel side-scan images are close to each other, see Fig. 2.
In contrast, methods [33][34][35][36] based on nearest neighbor fields (NNF) are more suitable for our problem, as they are able to perform efficient global search for best match directly on full image resolution. Nevertheless, NNF-based methods usually generate many outliers that are difficult to identify, due to lack of regularization. In this paper, we propose an effective dense matching method for side-scan image, by carefully modifying on a NNF-based seminal work called PatchMatch [7] to combine both dead-reckoning and imaging information to search for dense correspondences. To mitigate the influence of outliers to pose estimation, a robust RanSaC-based pose estimation pipeline is introduced.
@@ III-A1 Canonical Transformation
Data association across different SSS images is challenging, as the same physical area on the seabed will appear differently in the SSS images collected from different distances and angles. Furthermore, an SSS image will be distorted compared to an orthographic projection of the ensonified region and will vary in intensity as a function of the sonar position. To reduce such distortions, we apply a method [37] capable of transforming SSS images from different survey lines into a canonical representation, which includes two steps: intensity correction and sensor-independent slant range correction.
Intensity correction is performed to normalize the intensities across images based on the “Lambertian” model [38][39], with the SSS backscatter intensity being proportional to the cosine square [40] of the incidence angle. The slant range correction adjusts for the varying projection of side-scan pixels on the assumed horizontal seafloor. An adjustment of sine of the incidence angle is performed so that the pixels can be of fixed size in terms of horizontal range rather than slant range. More details can be found in [37].
@@ III-A2 Geo-referenced Image
To approximately check whether any two of the SSS images are overlapping, we follow the method in [26] and utilize the dead-reckoning data to geo-reference each pixel in the SSS images. This provides an approximate location of each pixel of an SSS image in global reference coordinates. These geo-references are also used to narrow the searching area when performing dense matching in data association.
@@ III-B Data Association
Once an overlapping area is found between a new SSS image and any image in the database, a data association process is performed to find dense pixel correspondences between the overlapping images. In this work, we propose to accomplish this by formulating it as a - (NNF) estimation problem, and solving it with computing the approximate NNF based on our modified version of PatchMatch [7], which is carefully tailored for side-scan images.
Specifically, define a nearest-neighbor field over all possible patch coordinates (locations of patch centers) in a side-scan image , so that given a patch coordinate in image , its corresponding nearest neighbor in an overlapping side-scan image can be denoted as,
The values of are referred to as and stored as a matrix with the same dimension of the image size of . Following the pipeline in [7], the estimation of has three main steps: initialization, random search and propagation, see Fig. 4 and Algorithm 1.
@@ III-B2 Random Search
Based on the initialized offset at each position, a random search is performed around it within a , which can be determined by prior knowledge of navigation drift and imaging resolution. Then, given a randomly selected position, we compare the patch distance against its current position. If the randomly selected position yields a smaller patch distance than the current position, then the offset is updated to represent the newly selected position. The distance function we use in this paper is Zero-mean Normalized Cross-Correlation (ZNCC) distance, which is invariant to linear brightness and contrast variations that often appear close to the nadir and the sides in side-scan images.
@@ III-B3 Propagation
Assuming that the patch offsets are likely to be the same locally, the propagation step aims to improve the values of using the known surrounding offsets. Specifically, let be the patch distance between the patch at in image and the patch at in , we take the new value for to be the minimum patch distance within the neighbors:
The effect is that if has a correct mapping and is in a coherent region, then all neighbors around will be filled with the correct mapping. However, side-scan images exhibit discontinuities due to multiple reasons, including canonical transformation induced centre-column discontinuity, noisy areas induced by AUV turning motion, as well as sensor artifacts. These continuities are challenging for the propagation. To address this, we mask out these areas and exclude them from the matching process.
@@ III-B4 Improving Data Association using Optimized Poses
The presets of patch size and max offset cannot be too high, as that would in general lower the patch matching accuracy, which makes it difficult to handle large offsets/drifts in cases where the paired images are distant from each other. To address this, we propose to improve the dense matching accuracy in distant paired images with optimized poses in an iterative fashion (see Fig. 3), which in turn helps to improve the pose accuracy. Here, a new parameter representing the iteration number of this process is introduced. Detailed discussion can be found in Section IV-C2.
@@ III-C Subframe Pose Estimation
Now we show how each pair of the pixel correspondences is used to construct a measurement model. Based on that, we propose a robust pose estimation pipeline, which is able to estimate the relative pose between a pair of subframes of SSS images, by iteratively identifying a set of good quality inliers from dense correspondences using RanSaC. In the following, we assume that each side-scan image has been evenly divided into multiple subframes and they are used as unit image for estimation instead of the whole one, see Fig. 5 left.
@@ III-C2 Relative Pose Estimation
Given a set of pixel correspondences between two overlapped subframes and , the centre ping poses of both subframes and , and the observed D landmarks = can be estimated via minimizing the minus log of least squares cost, assuming the measurement noises as Gaussian:
with the first term on the right side being the side-scan measurement cost calculated by Eq. 3, the second term the odometry measurement cost and the third term a prior cost. Specifically, is the relative pose between the centre pings of the subframes, and is the measurement obtained from dead-reckoning data. The odometry term is crucial as it helps to relieve the underconstrained issue in side-scan measurement and ensure convergence to a desired minimum. is a prior model that sets one of the poses (e.g., here) fixed and only adjust for the other, which is reasonable as we aims to estimate the relative pose between them rather than estimate both.
is the covariance matrix for each side-scan bin measurement defined in Eq. 3. The variance in the range measurement is dominated by the discretization of the range in the side-scan, varying from - depending on the resolution of the image. For the second row restricting the landmark to lie in the plane, the variance should grow as the square of the horizontal distance from the sonar. Since the sonar is mostly used at shallow grazing angles, we can instead use the range so that:
where is the beam width in radians. is the odometry covariance and is set proportional to the distance between the two dead-reckoning poses. A factor graph describing the costs in Eq. 5 as factors, is illustrated in Fig. 5 bottom right (dashed rectangle).
@@
The core idea of this pipeline is the consideration of both measurement cost ( and ) in the unsampled data, and in the sampled data. The former measures how well the hypothesis model fits the unsampled data, while the latter indicates how well the optimization converges using the sampled data. We found experimentally that only using the measurement cost to determine the best model could be biased in many cases, while combining with optimization cost helps mitigate this issue by avoiding bad convergence. Therefore, we only update the current best model when both the costs are decreased. In additions, the final cost values also serve as indicators against empirical-based thresholds (see Section IV-4), to filter out the poorly-estimated relative pose from being used as loop closure edge in pose graph optimization.